OpenStudy (anonymous):

Find the derivative y=x^2 * cosx y'= Use y'=udv + vdu

zepdrix (zepdrix):

Here is the setup for product rule. \[\large y'=\color{royalblue}{(x^2)'}\cos x+x^2\color{royalblue}{(\cos x)'}\] We need to take the derivative of the blue terms. That is what the primes indicate. Understand the setup based on the rule they told you to use?

OpenStudy (anonymous):

Yes I do!

zepdrix (zepdrix):

Understand how to finish it? :) Gotta remember your trig derivatives!! :D

OpenStudy (anonymous):

Don't you have to simplify it?

zepdrix (zepdrix):

You have to `take the derivative` now. All we did was setup the product rule so far.

zepdrix (zepdrix):

Take the derivative of the blue terms. Those are the ones that need work still.

OpenStudy (anonymous):

I'm afraid I'm a bit confused on finding the derivative :/

zepdrix (zepdrix):

So we've successfully setup the product rule. \[\large y'=\color{royalblue}{(x^2)'}\cos x+x^2\color{royalblue}{(\cos x)'}\] Let's take it piece by piece. We first want the derivative of x^2. We can simply apply the power rule to find it. Member the power rule? :O

zepdrix (zepdrix):

Oooo is your picture a watermelon? XD Oh man that's fancy!

OpenStudy (anonymous):

Refresh my memory please >_< sounds awfully familiar!

OpenStudy (anonymous):

And yes it is!! :D I struggle just to cut one normally at times XD

zepdrix (zepdrix):

Power Rule for Derivatives: ~Bring the power down as a coefficient in front. ~Decrease the power by 1. Example:\[\large (x^3)' \qquad = \qquad 3x^{3-1} \qquad = \qquad 3x^2\]

OpenStudy (anonymous):

OH YES. I remember it now!

zepdrix (zepdrix):

So what would the derivative of \(\large x^2\) be? :D

OpenStudy (anonymous):

2x...?

zepdrix (zepdrix):

\[\large y'=\color{orangered}{(2x)}\cos x+x^2\color{royalblue}{(\cos x)'}\] Good. So we've found the derivative of that first blue term. The second term, there really isn't much of a process that we'll want to go through. You just need to memorize the derivative of cosine x.

OpenStudy (anonymous):

Alrighty! So what's next?

zepdrix (zepdrix):

What's the derivative of \(\large \cos x\) ? Look it up in your notes somewhere if you have to! XD

OpenStudy (anonymous):

Goodness...is it 0 or something? lol

zepdrix (zepdrix):

noooooooooooo :O

OpenStudy (anonymous):

Oh nooo DX my notes are all messy.

zepdrix (zepdrix):

http://en.wikipedia.org/wiki/Differentiation_of_trigonometric_functions Look at the chart in the upper right, it should be on that list.

zepdrix (zepdrix):

Our `function` is \(\large \cos(x)\), what should the `derivative` be?

OpenStudy (anonymous):

Ahaa so that's what my notes said. -sin(x)?

zepdrix (zepdrix):

Yayyy you found it :D \[\large y'=\color{orangered}{(2x)}\cos x+x^2\color{orangered}{(-\sin x)}\]

zepdrix (zepdrix):

One final step, to clean it up might be to move the negative sign in front as subtraction. It will just look a little nicer.

zepdrix (zepdrix):

\[\large y'=2x \cos x-x^2 \sin x\]

OpenStudy (anonymous):

Awesome :D so that is it?

zepdrix (zepdrix):

mhm -_-